OpenStudy (anonymous):

Fan and Medal What number should be added to the expression x^2 + 10x to change it into a perfect square trinomial? A. 5 B. 10 C. 25 D. 100

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

25 :)

OpenStudy (anonymous):

x^2+2*5x+25 = x^2+2*5x+5^2 = (x+5)^2